OpenStudy (anonymous):

25% of what number is 6.25

OpenStudy (agl202):

What do you think?

sammixboo (sammixboo):

Once again :P use this formula \(\rm \color{blueviolet}{\dfrac{part}{whole}=\dfrac{percent}{100}}\) Where 25 is the percent and 6.25 is the part Plug in the values \(\rm \color{blueviolet}{\dfrac{6.25}{whole}=\dfrac{25}{100}}\) Do you know how to solve this?

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

multiply 6.25 by 4 and that's the answer

OpenStudy (agl202):

yes that's is correct.

OpenStudy (anonymous):

wow thanks yah its 25

sammixboo (sammixboo):

To solve \(\rm \color{blueviolet}{\dfrac{6.25}{whole}=\dfrac{25}{100}}\) you have to cross multiply 6.25 by 100  So what is 6.25 x 100?